UT 01 Oct 2021 - Split Night - PEPSI/PETS - D/E




Night Plan

  • PEPSI completes at 00:00hr - (UT 00:00)
  • Reconfigure to LUCI-Bino - start at 00:15 AZ time
  • LUCI-2 Checkout (highest priority)
    • Scripts located in ~lbto/Service_Observing/Restart_2021/LUCI/LUCI_2021_Restart_V1.6X - last version of OT is 1.63 (as of DPM meeting)
      • Rot ZP already exported in RotZP sub-directory - suggest doing this and AGW transform as LUCI-2 only to save time
    • Rotator ZP w/MOS mask - 0.5hr - seeing 1.2" or better
      • Stone P sets at 02hr
      • Stone A sets at 04hr
      • Do both masks but note Stone P has some outliers, #36 is the worst. If Stone P is done, also do Stone A. Stone A had no outliers
    • AGW Transform - 1-1.5hr - seeing 1.2" or better
      • N30 fieldstop w/N375 camera, H-filter, flexure comp = off (DIT=2.51, but DIT=4 is also OK)
      • Stone_A3_LUCI.list
      • Stone_P1.list
    • Imaging dither test/Science - Bino
      • GRB 200826A - J-band bino 1.25 hr - seeing 1.2" or better - up all night
        • OT file located in Science_Dither_Test/ Grb200826A _OT-LUCIonly_V3.xml
        • LUCI-1 and LUCI-2 versions in OT file
      • NGC 40 - visible all night - any seeing, any conditions
        • can be generated from LUCI_2021_Restart_V1.6X.xml
    • NS LS dither test - Bino
      • Not exported yet from LUCI_2021_Restart_V1.6X.xml
      • Heracles 5143 - G200 longslit - any conditions - any seeing 1hr (either Heracles or 5143 works for NSIGUI)
      • Uranus or 799 - G210 longslit - any conditions - any seeing 1hr
    • Photometric Standards - Bino - any seeing - clear
      • Not exported yet from LUCI_2021_Restart_V1.6X.xml
  • Poor Weather Backup
    • Execute all that can be done for LUCI-2 checks then move to poor weather (including some LUCI-Bino Standards - clear weather)
    • LBC Active TMS - 3hrs - any conditions
    • LBC Standards - clear any seeing
    • MODS Standards - clear any seeing


Summary

High humidity lingered through most of the night, but we were able to open briefly at the end of the first half and then again around 18-deg morning twilight. We executed a few dithers on-off source on NGC40 and we then acquired a long-slit target (Q0636+680) and dithered up/down by +30/-30" to see how well the object remained centered. This was not too conclusive as only one down/up cycle could be done before it got too light, and while the down position was centered, the up one was offset to the left of the 1" slit. Seeing was about 1".

One concern was collimation on SX --- we opened briefly at the end of the first half and tried to collimate with PEPSI but there was a lot of coma on SX and it did not converge by handover so we gave up and reconfigured to LUCI. The PEPSI target was at low elevation when we tried to collimate. The first LUCI target was also at low elevation and the SX side did not seem to converge (but must have dipped below the threshold since observations started). This target was in the N. We tried another target, also in the N, and collimation was a little better but it was again slow to converge. If there were more time, I would have tried a target somewhere else in the sky, because this leaves some doubt as to the collimation on SX - whether there was a problem or it was just dome seeing. L1 was already checked on sky, so if nothing has changed, then I would imagine that there is not a problem and what we were seeing was an environmental effect (??)

Night Log

04:53 The RH finally dropped and David is preparing to open. The PEPSI transit is still viable. LUCI SW and LBCs are run up. In L2, I tweaked the FM4 position via the AFC tab to shift the field by dx ~ +15 pixels.

05:11 RH shot up and is now 99.9%.

Note re: IT 8470 (PEPSI-PFU transit observation during which the star was lost several times on the DX side and the series of guider images showed periodic "flux flashing" (saturation of star in multiple pixels one every n-th image, whereas the peak pixel had ~30,000 counts on other images). Tonight's transit observation is of the same target and Ilya had a test to make - details unknown and test so far not done because we have not opened.

06:32 David is opening. Too late for the transit, but Ilya is doing a preset to the star to see if the guider grabs the same/correct star. Both guiders failed - "ERROR: full Guide image didn't arrive in time". David is restarting GCS. Earlier Ilya had restarted/rebooted both AZCAMS - this may explain it, as the initial test preset that David did worked fine.

  • There is not a button to reconnect GCS to Azcam - one has to start GCS after rebooting AZcams.

The test preset was made to try to reproduce the problem when this target was observed a couple of weeks ago - at that time, GCS selected the wrong star on one side.

07:10 The RH rose again and David closed.

07:05 David authorized for LUCI.

11:51 David is opening; RH has been below the limit for 30-mn. There are clouds.

12:00 Slew to NGC40 for imaging dither test. It is low, 35 deg - David is making the pointing adjustments. Collimated on a nearby star.

12:19 NGC40 imaging OB started. Collimation on SX is not good. It wasn't good with PEPSI either, but we stopped and reconfigured before allowing it to finish. On SX: WFE ~ 540-700 and FWHM 1.8" but on DX: WFE ~ 200 and FWHM 1.0" (see screenshot below).

L1: 2 and L2: 4 are on target L1: 3 and L2: 5 are at sky pos'n and not guided

We ran through a few dither positions: L1: 2-6 and L2: 4-8; before going to a new target to dither up/down in the slit.

12:31 Slew to Q0636+680 and acquire in 1" slit. Dither up/down. This will help us see whether the poor SX collimation is a more general problem (or was it dome seeing or something related to the low elevation that both PEPSI target & NGC40 had) and will check for misalignment between slit & detxy Y.

elevation: 55 deg, pointing to N and wind is from NW. The SX side collimation is not converging very quickly. Oddly, LUCI images started. WFE may have dipped below threshold and then rose... I wasn't watching it that carefully. SX: WFE ~1000 and not converging. FWHM ~1.2" but finally WFE down to 261. DX: WFE ~ 600 and FWHM ~ 1"

Observer error (forgot to click "set" for the slit position on L1) and as a result commanded offsets of 120" on SX and hit the copointing limit. Offsets were not made and I manually sent the correct offsets. L2 is really what we care about anyway, not L1.

On confirmatory thru-slit image luci2.20211001.0012.fits, the object looks well-centered in the 1" slit (though overfills it). Will execute a few dithers in imaging mode: +/- 30" before it gets too light. On L2: 13, DETXY -30" looks centered On L2: 14, DETXY +30" looks offset to the left edge by about 1.5 pixels (N375).

By-eye centroids for the star in both positions and the slit center at both positions suggest a 0.15 deg offset between slit and detxy y-axis, but this is pretty tentative.

12:54 David is closing up. It is too light to continue guiding.

12:50: Lost WFS on SX side - image is partially black/shadowed. OK on DX side.

-- BarryRothberg - 29 Sep 2021
  • Screenshot from 2021-10-01 05-24-39.png:
    Screenshot from 2021-10-01 05-24-39.png
Topic revision: r7 - 01 Oct 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback